# Reasoning Limitations of Multimodal Large Language Models. A case study of Bongard Problems

### v1

@@ Abstract
Abstract visual reasoning (AVR) encompasses a suite of tasks whose solving requires the ability to discover common concepts underlying the set of pictures through an analogy-making process, similarly to human IQ tests. Bongard Problems (BPs), proposed in 1968, constitute a fundamental challenge in this domain mainly due to their requirement to combine visual reasoning and verbal description. This work poses a question whether multimodal large language models (MLLMs) inherently designed to combine vision and language are capable of tackling BPs. To this end, we propose a set of diverse MLLM-suited strategies to tackle BPs and examine four popular proprietary MLLMs: GPT-4o, GPT-4 Turbo, Gemini 1.5 Pro, and Claude 3.5 Sonnet, and four open models: InternVL2-8B, LLaVa-1.6 Mistral-7B, Phi-3.5-Vision, and Pixtral 12B. The above MLLMs are compared on three BP datasets: a set of original BP instances relying on synthetic, geometry-based images and two recent datasets based on real-world images, i.e., Bongard-HOI and Bongard-OpenWorld. The experiments reveal significant limitations of MLLMs in solving BPs. In particular, the models struggle to solve the classical set of synthetic BPs, despite their visual simplicity. Though their performance ameliorates on real-world concepts expressed in Bongard-HOI and Bongard-OpenWorld, the models still have difficulty in utilizing new information to improve their predictions, as well as utilizing a dialog context window effectively. To capture the reasons of performance discrepancy between synthetic and real-world AVR domains, we propose Bongard-RWR, a new BP dataset consisting of real-world images that translates concepts from hand-crafted synthetic BPs to real-world concepts. The MLLMs’ results on Bongard-RWR suggest that their poor performance on classical BPs is not due to domain specificity but rather reflects their general AVR limitations.
(a) Bongard Problem #31 (Bongard, 1970)
(b) Bongard-RWR (This work)
(c) Bongard HOI (Jiang et al., 2022)
(d) Bongard-OpenWorld (Wu et al., 2024)
Figure 1: Bongard Problems. In this work, we consider BPs with both synthetic and real-world images. (a) A manually designed BP #31. Left: One line. Right: Two lines. (b) A real-world representation of BP #31 from the proposed Bongard-RWR dataset introduced in Section 4. Left: One line. Right: Two lines. (c) Bongard HOI: Left: A person jumping on a surfboard. Right: Not a person jumping on a surfboard. (d) Bongard-OpenWorld: Left: An abstract painting. Right: Not an abstract painting. Descriptions of the Left / Right sides come from the respective datasets.
@@ 1 Introduction
Multiple problems that necessitate combined perception and reasoning have been identified (Hofstadter, 1999). Among these tasks are Bongard Problems (BPs), introduced by Bongard (1968; 1970). Initial BPs were designed manually, leading to the formulation of a few hundred task instances by individual contributors (Foundalis, 2006b). A typical BP consists of two sides, left and right, each comprising six image panels arranged in a grid. All images on one side illustrate a shared concept absent in the images on the opposite side. The task is to identify the underlying rule that differentiates the sides and articulate it in natural language. Initial BPs (Bongard, 1968), akin to human IQ tests, featured abstract 2D geometric shapes, putting the focus on abstract reasoning. However, recent works have expanded the set of BPs to include real-world images, which broadens the scope of presented objects, attributes and relations. Specifically, the matrices in Bongard HOI (Jiang et al., 2022) depict human-object interactions, while Bongard-OpenWorld (Wu et al., 2024) employs open-world free-form concepts, increasing the diversity of featured scenes. Figs. 1(a)1(c)1(d) illustrate examples of problems from the three above-mentioned datasets.
A central theme in BPs is recognition of concepts in a context-dependent manner, as object representations need to be formed specifically for the presented matrix, rather than described a priori (Linhares, 2000). For example, consider the matrix in Fig. 1(a) – an analysis restricted to its left side may yield multiple concepts, such as the presence of curves or an object centered in the image. Only through a comprehensive understanding of both matrix sides one can recognize that the left side depicts a single line, while the right side presents two lines. Such concept-based tasks were argued to promote a more accurate evaluation of a system’s generalization ability and its capacity for abstraction (Mitchell, 2021; Odouard & Mitchell, 2022). Moreover, the concepts in BPs are illustrated with several image examples, which positions the task within a few-shot learning setting (Fei-Fei et al., 2006; Wang et al., 2020). In contrast to other abstract reasoning problems, such as Raven’s Progressive Matrices (RPMs) (Raven, 1936; Raven & Court, 1998; Małkiński & Mańdziuk, 2022) that have recently witnessed the development of large-scale benchmarks (Barrett et al., 2018; Zhang et al., 2019), BPs allow to assess system’s ability to derive concepts from a limited set of examples (typically six images per matrix side). The above aspects make BPs a valuable testbed for assessing abstract reasoning abilities of AI models.
@@
The quest to build systems capable of forming abstract concepts dates back to the 1950s (McCarthy et al., 2006). The advent of Deep Learning (DL) opened new possibilities to tackle BPs (Kharagorgiev, 2018; Nie et al., 2020). However, despite significant advancements, methods for consistently solving BPs (and other problems that involve abstract reasoning) are still lacking (Mitchell, 2021; van der Maas et al., 2021; Stabinger et al., 2021). Typically, DL approaches omit the generation of natural language answers by casting BP into a binary classification task, in which a test image had to be assigned to the matching side of the matrix. Conversely, a parallel stream of research on large language models (LLMs) demonstrated promising results in open-ended language generation (Brown et al., 2020). In particular, LLMs were applied to selected AVR tasks (Webb et al., 2023), though, lately Xu et al. (2024a) pointed certain LLM limitations in solving AVR problems represented as text despite using information lossless translation through direct-grid encoding. Recent works have combined the vision and language modalities into multimodal large language models (MLLMs) (Achiam et al., 2023; Reid et al., 2024; Anthropic, 2024), inviting their application to diverse tasks (Yin et al., 2023; Wu et al., 2023). Motivated by these recent developments we examine the reasoning capabilities of MLLMs in solving BPs.
Contributions.
The main contribution of this paper is four-fold.
(1) For the first time in the literature, we consider BPs in the context of MLLMs and propose a diverse set of strategies to solve BP instances in two setups: open-ended language generation and binary classification.
@@
(3) To further examine the main difficulties faced by MLLMs in solving both types of BPs (synthetic and real world ones) we introduce a focused dataset of BPs (Bongard-RWR) comprising real-world images that represent concepts from synthetic BPs using real world images. Thanks to relying on the same abstract concepts as synthetic BPs, Bongard-RWR facilitates direct comparisons of the MLLMs performance in both domains.
@@ 2 Related Work
The AVR field encompasses a broad set of problems aimed at studying various aspects of visual cognition (Gardner & Richards, 2006; Małkiński & Mańdziuk, 2023). Recent DL research in this domain gravitated towards utilizing certain well-established datasets, e.g. with visual analogies (Hill et al., 2019; Webb et al., 2020) or RPMs (Zhang et al., 2019; Barrett et al., 2018), to measure the progress of DL models. However, such benchmarks evaluate system performance in learning a particular task, rather than assessing its general ability to acquire new AVR skills. To address this limitation, certain tasks have adopted few-shot learning setups, requiring models to learn from a few demonstrations, as exemplified by SVRT (Fleuret et al., 2011) or Bongard-LOGO (Nie et al., 2020). Nonetheless, these benchmarks follow a discriminative setting where a set of possible answers is provided. Conversely, other datasets such as ARC (Chollet, 2019) or PQA (Qi et al., 2021) pose a generative challenge, which may be considered more difficult due to its open-ended nature. In addition to synthetic tasks featuring 2D geometric shapes, certain datasets present analogous reasoning tasks using real-world images (Teney et al., 2020; Ichien et al., 2021; Bitton et al., 2023). This approach extends the range of concepts that can be expressed and, above all, allows employing models pre-trained on large image datasets. In this work, we concentrate on several BP datasets that present a few-shot learning challenge, cover both synthetic and real-world images, and consider settings involving both binary classification and answer generation in natural language.
@@
Initial approaches to tackle BPs involved cognitive architectures (Foundalis, 2006a), program synthesis coupled with inductive logic programming (Saito & Nakano, 1996; Sonwane et al., 2021), and the application of Bayesian inference within a visual language framework (Depeweg et al., 2018; 2024). Kharagorgiev (2018) trained a convolutional network on a generated synthetic dataset with geometric shapes and applied a one-level decision tree to solve BPs framed as a binary classification task. Nie et al. (2020) introduced Bongard-LOGO with synthetically generated BPs and used it to evaluate CNN-based models focused on meta-learning (Snell et al., 2017; Mishra et al., 2018; Lee et al., 2019; Raghu et al., 2020; Chen et al., 2021) and relational reasoning (Barrett et al., 2018). Jiang et al. (2022) applied the Relation Network (Santoro et al., 2017) to objects detected with Faster R-CNN (Ren et al., 2015) and employed the model to solve matrices from the real world Bongard HOI dataset. Despite high diversity of approaches, none of them has fully addressed the abstract and open-ended nature of BPs. Most related to our work, Wu et al. (2024) considered hybrid approaches that caption each image panel with an image-to-text model and applied LLMs for processing these text descriptions. Differently, in this work we focus on MLLMs that are inherently capable of jointly processing images and text.
@@
MLLMs haven’t been yet applied to tackle BPs, though they were applied to several related tasks. Initial works focused on LLMs and evaluating their abstract reasoning performance in simplified analogy tasks. Webb et al. (2023) showed that GPT-3 and GPT-4 (text-only variants) performed on the human level, or even outcompeted humans, in certain RPM-like tasks in a zero-shot manner without additional fine-tuning. However, they represented the image objects as text using a fixed small vocabulary, thus omitting the need for identifying concepts from open-ended shapes, a key challenge of BPs. Recent research concerning the evaluation of abstract reasoning skills of LLMs concentrates around the Abstraction and Reasoning Corpus (ARC) task (Chollet, 2019). It was demonstrated that LLMs can solve certain ARC problems transformed to the text domain (Moskvichev et al., 2023; Mirchandani et al., 2023; Camposampiero et al., 2023; Xu et al., 2024b). Despite these important stepping stones, the text-based representation taken in these works simplifies the perception task by presenting the model with pre-existing higher level representations. Only recently, thanks to the appearance of MLLMs, vision and text started to be treated jointly in a unified manner. Cao et al. (2024) proposed a suite of AVR tasks to compare MLLM and human performance. Jiang et al. (2024) assessed AVR skills of MLLMs on an introduced multidimensional benchmark combining AVR and perceptual questions. Our work complements this stream of research by exploring BPs, a fundamental task in the field, and providing insights into MLLM analogy-making performance in synthetic and real-world domains.
@@ 3 Solving BPs with MLLMs
3.1 Prompting Strategies for Natural Language Answer Generation
We start by defining the strategies for generating answers in natural language. In each strategy, the model receives a general description of Bongard Problem with two BP examples with correct answers. Additionally, besides this generic introductory information, a given task ${\cal BP}^{X}$ to be solved is presented in a strategy-specific way. Appendix A.4 presents the exact prompt formulations.
@@ 3.1 Prompting Strategies for Natural Language Answer Generation
Descriptive (Fig. 2b). Defines a more granular approach in which the model is first requested to generate a textual description of each image panel of the matrix. Each description is generated in a separate context, such that the model doesn’t have access to the prior panels nor to their descriptions. Next, the model is requested to provide an answer to the problem based only on the generated textual descriptions of all image panels.
@@
Contrastive-direct (Fig. 2d with a dashed element). Extends Contrastive by including the image of the whole matrix together with textual descriptions of differences within each panel pair.
@@
3.2 Evaluation of Solutions Expressed in Natural Language
The correct answer to a BP may be formulated in natural language in many different ways. To account for this inherent variability, we utilize a model-based approach to assess whether the generated answer $\widehat{y}$ matches the ground-truth $y$. In the proposed setting, an MLLM ensemble receives both $\widehat{y}$ and $y$ and is requested to output a binary yes/no answer whether both descriptions refer to the same concept (see Fig. 3a). Specifically, we assess the correctness of $\widehat{y}$ using all four considered proprietary MLLMs and count it as correct if at least two models agree with this class (see details in Appendix B). In contrast to solving the BP task, which requires abstract reasoning abilities, the task of determining whether two answers refer to the same concept boils down to assessing the semantic similarity between two texts, and MLLMs are known to excel in such setup (Lu et al., 2024).
@@ 3.3 Binary classification formulations
To dive deeper into the AVR capabilities of the studied models, we cast the BP task into three binary classification settings, reducing the task’s difficulty. Firstly, we provide the model with the image of the whole matrix ${\cal BP}^{X}$ along with a possible solution, and the model is prompted to generate a binary score assessing the correctness of the provided answer (Fig. 3b). Two settings are considered, in which the solution is formed by either the actual ground-truth answer (the expected answer is yes), or by an incorrect answer taken from a different BP matrix (the expected answer is no). Secondly, we follow a setting exemplified in the Bongard-LOGO dataset (Nie et al., 2020) in which two test images have to be classified to different sides of the problem (Fig. 3c). To this end, we take two test images corresponding to the respective sides of the matrix, randomly shuffle the images, and request the model to determine the side to which each image belongs. In synthetic BPs we create the test set by removing the 6th image from each side of the matrix, while in BPs from Bongard HOI, Bongard-OpenWorld and Bongard-RWR we use the additional test images. We refer to these three formulations as Ground-truth, Incorrect Label, and Images to Sides (see prompts in Appendix A.3).
4 Bongard-RWR: Synthetic BPs expressed in real-world images

### v2

# A Case Study of Bongard Problems

@@ Abstract
Abstract visual reasoning (AVR) involves discovering shared concepts across images through analogy, akin to solving IQ test problems. Bongard Problems (BPs) remain a key challenge in AVR, requiring both visual reasoning and verbal description. We investigate whether multimodal large language models (MLLMs) can solve BPs by formulating a set of diverse MLLM-suited solution strategies and testing $4$ proprietary and $4$ open-access models on $3$ BP datasets featuring synthetic (classic BPs) and real-world (Bongard HOI and Bongard-OpenWorld) images. Despite some successes on real-world datasets, MLLMs struggle with synthetic BPs. To explore this gap, we introduce Bongard-RWR, a dataset representing synthetic BP concepts using real-world images. Our findings suggest that weak MLLM performance on classical BPs is not due to the domain specificity, but rather comes from their general AVR limitations. Code and dataset are available at: https://github.com/pavonism/bongard-rwr
Multimodal Large Language Models, Abstract Visual Reasoning, Bongard Problems
@@ 1 Introduction
(a) Bongard Problem #$\#31$
(b) Bongard-RWR #$\#31$
(c) Bongard Problem #$\#36$
(d) Bongard-RWR #$\#36$
Figure 1: Bongard Problems. (a), (c): Manually-designed synthetic BPs (Bongard, 1970). (b), (d): Their real-world equivalents from the Bongard-RWR dataset proposed in this work. BP #$\#31$: Left: One line. Right: Two lines. BP #$\#36$: Left: Triangle above circle. Right: Circle above triangle.
Multiple problems that necessitate combined perception and reasoning have been identified (Hofstadter, 1999). Among these tasks are Bongard Problems (BPs), introduced by Bongard (1968; 1970). Initial BPs were designed manually, leading to the formulation of a few hundred task instances by individual contributors (Foundalis, 2006b). A typical BP consists of two sides, left and right, each comprising six image panels arranged in a grid (Figs. 1(a)1(c)). All images on one side illustrate a shared concept absent in the images on the opposite side. The task is to identify the underlying rule that differentiates the sides and articulate it in natural language. Initial BPs (Bongard, 1968), akin to human IQ tests, featured abstract 2D geometric shapes, putting the focus on abstract reasoning. However, recent works have expanded the set of BPs to include real-world images, which broadens the scope of presented objects, attributes and relations. Specifically, the matrices in Bongard HOI (Jiang et al., 2022) depict human-object interactions, while Bongard-OpenWorld (Wu et al., 2024) employs open-world free-form concepts, increasing the diversity of featured scenes.
A central theme in BPs is recognition of concepts in a context-dependent manner, as object representations need to be formed specifically for the presented matrix, rather than described a priori (Linhares, 2000). For example, consider the matrix in Fig. 1(a) – an analysis restricted to its left side may yield multiple concepts, such as the presence of curves or an object centered in the image. Only through a comprehensive understanding of both matrix sides one can recognize that the left side depicts a single line, while the right side presents two lines. Such concept-based tasks are argued to offer a more accurate evaluation of the model’s generalization ability and its capacity for abstraction (Mitchell, 2021; Odouard & Mitchell, 2022). In contrast to other abstract reasoning problems, such as Raven’s Progressive Matrices (RPMs) (Raven, 1936; Raven & Court, 1998; Małkiński & Mańdziuk, 2025a) that have recently witnessed the development of large-scale benchmarks (Barrett et al., 2018; Zhang et al., 2019; Małkiński & Mańdziuk, 2025c), BPs allow to assess model’s ability to derive concepts from a limited set of examples (typically six images per matrix side), which positions the task within a few-shot learning setting (Fei-Fei et al., 2006; Wang et al., 2020). The above aspects make BPs a valuable testbed for assessing abstract reasoning abilities of AI models.
@@
The quest to build systems capable of forming abstract concepts dates back to the 1950s (McCarthy et al., 2006). The advent of Deep Learning (DL) opened new possibilities to tackle BPs (Kharagorgiev, 2018; Nie et al., 2020). However, despite significant advancements, methods for consistently solving BPs (and other problems that involve abstract reasoning) are still lacking (Mitchell, 2021; van der Maas et al., 2021; Stabinger et al., 2021). Typically, DL approaches omit the generation of natural language answers by casting BP into a binary classification task, in which a test image had to be assigned to the matching side of the matrix. Conversely, a parallel stream of research on large language models (LLMs) demonstrated promising results in open-ended language generation (Brown et al., 2020). In particular, LLMs were applied to selected AVR tasks (Webb et al., 2023), though, lately Xu et al. (2024) pointed certain LLM limitations in solving AVR problems represented as text despite using information lossless translation through direct-grid encoding. Recent works have combined the vision and language modalities into multimodal large language models (MLLMs) (Achiam et al., 2023; Reid et al., 2024; Anthropic, 2024), inviting their application to diverse tasks (Yin et al., 2024; Wu et al., 2023). Motivated by these recent developments we examine the reasoning capabilities of MLLMs in solving BPs.
Contribution.
The contribution of this paper is fourfold.
(1) We consider BPs in the context of MLLMs and propose a diverse set of strategies to solve BP instances in two setups: open-ended language generation and binary classification.
@@
(3) To further examine the observed MLLM limitations in solving both synthetic and real-world BPs, we introduce a focused dataset of BPs (Bongard-RWR, Figs. 1(b)1(d)) that represent concepts from synthetic BPs using real world images. Thanks to relying on the same abstract concepts as synthetic BPs, Bongard-RWR facilitates direct comparisons of the MLLM performance in both domains.
@@ 2 Related Work
The AVR field encompasses a broad set of problems aimed at studying various aspects of visual cognition (Gardner & Richards, 2006; Małkiński & Mańdziuk, 2023). Recent DL research in this domain gravitated towards utilizing certain well-established datasets, e.g. with visual analogies (Hill et al., 2019; Webb et al., 2020) or RPMs (Zhang et al., 2019; Barrett et al., 2018), to measure the progress of DL models (Małkiński & Mańdziuk, 2024a, b, 2025b). However, such benchmarks evaluate system performance in learning a particular task, rather than assessing its general ability to acquire new AVR skills. To address this limitation, certain tasks have adopted few-shot learning setups, requiring models to learn from a few demonstrations, as exemplified by SVRT (Fleuret et al., 2011) or Bongard-LOGO (Nie et al., 2020). Nonetheless, these benchmarks follow a discriminative setting where a set of possible answers is provided. Conversely, other datasets such as ARC (Chollet, 2019) or PQA (Qi et al., 2021) pose a generative challenge, which may be considered more difficult due to its open-ended nature. In addition to synthetic tasks featuring 2D geometric shapes, certain datasets present analogous reasoning tasks using real-world images (Teney et al., 2020; Ichien et al., 2021; Bitton et al., 2023). This approach extends the range of concepts that can be expressed and, above all, allows employing models pre-trained on large image datasets. In this work, we concentrate on several BP datasets that present a few-shot learning challenge, cover both synthetic and real-world images, and consider settings involving both binary classification and answer generation in natural language.
@@
Initial approaches to tackle BPs involved cognitive architectures (Foundalis, 2006a), program synthesis coupled with inductive logic programming (Saito & Nakano, 1996; Sonwane et al., 2021), and the application of Bayesian inference within a visual language framework (Depeweg et al., 2018, 2024). (Kharagorgiev, 2018) trained a convolutional network on a generated synthetic dataset with geometric shapes and applied a one-level decision tree to solve BPs framed as a binary classification task. Nie et al. (2020) introduced Bongard-LOGO with synthetically generated BPs and used it to evaluate CNN-based models focused on meta-learning (Snell et al., 2017; Mishra et al., 2018; Lee et al., 2019; Raghu et al., 2020; Chen et al., 2021) and relational reasoning (Barrett et al., 2018). Jiang et al. (2022) applied the Relation Network (Santoro et al., 2017) to objects detected with Faster R-CNN (Ren et al., 2015) and employed the model to solve matrices from the real world Bongard HOI dataset. Despite high diversity of approaches, none of them has fully addressed the abstract and open-ended nature of BPs. Most related to our work, Wu et al. (2024) considered hybrid approaches that caption each image panel with an image-to-text model and applied LLMs for processing these text descriptions. Differently, in this work we focus on MLLMs that are inherently capable of jointly processing images and text.
@@
While the abstract reasoning capabilities of MLLMs have only recently begun to be explored, they have been studied more extensively in several related tasks. Initial works focused on LLMs and evaluating their abstract reasoning performance in simplified analogy tasks. Webb et al. (2023) showed that GPT-3 and GPT-4 (text-only variants) performed on the human level, or even outcompeted humans, in certain RPM-like tasks in a zero-shot manner without additional fine-tuning. However, they represented the image objects as text using a fixed small vocabulary, thus omitting the need for identifying concepts from open-ended shapes, a key challenge of BPs. Recent research concerning the evaluation of abstract reasoning skills of LLMs concentrates around the Abstraction and Reasoning Corpus (ARC) task (Chollet, 2019). It was demonstrated that LLMs can solve certain ARC problems transformed to the text domain (Moskvichev et al., 2023; Mirchandani et al., 2023; Camposampiero et al., 2023; Xu et al., 2024). Despite these important stepping stones, the text-based representation taken in these works simplifies the perception task by presenting the model with pre-existing higher level representations. Only recently, thanks to the appearance of MLLMs, vision and text started to be treated jointly in a unified manner. Perception Test (Patraucean et al., 2023) assessed multimodal video understanding across various skills and types of reasoning. Cao et al. (2024) proposed a suite of AVR tasks to compare MLLM and human performance. Jiang et al. (2024) assessed AVR skills of MLLMs on an introduced multidimensional benchmark combining AVR and perceptual questions. KiVA (Yiu et al., 2025) evaluated how models generalize visual transformations to analogical reasoning tasks inspired by the developmental psychology. Most relevant to our paper is the concurrent work of Wüst et al. (2024), which investigated the capabilities of MLLMs in solving synthetic BPs, using both open-ended text generation and multi-class concept classification setups. Our work complements this stream of research by exploring BPs and providing comparative insights into MLLM analogy-making performance in both synthetic and real-world domains.
@@ 3 Solving BPs with MLLMs
3.1 Answer Generation in Natural Language
We start by defining the strategies for generating answers in natural language. In each strategy, the model receives a general description of Bongard Problem with two BP examples with correct answers. Additionally, besides this generic introductory information, a given task ${\cal BP}^{X}$ to be solved is presented in a strategy-specific way. Appendix I.4 presents the exact prompt formulations.
@@ 3.1 Answer Generation in Natural Language
Descriptive (Fig. 2b). Defines a more granular approach in which the model is first requested to generate a textual description of each image panel of the matrix. Each description is generated in a separate context, such that the model doesn’t have access to the prior panels nor to their descriptions. Next, the model is requested to provide an answer to the problem based only on the generated textual descriptions of all image panels. In this strategy and in all subsequent ones, we evaluate only the final answer provided by the model, without considering the intermediate outputs.
@@
Contrastive-direct (Fig. 2d with a dashed element). Extends Contrastive by adding the image of the whole matrix to textual descriptions of differences within each panel pair.
@@
3.2 Evaluation of Natural Language Solutions
The correct answer to a BP may be formulated in natural language in many different ways. To account for this inherent variability, we utilize a model-based approach to assess whether the generated answer $\widehat{y}$ matches the ground-truth $y$. In the proposed setting, an MLLM ensemble receives both $\widehat{y}$ and $y$ and is requested to output a binary yes/no answer whether both descriptions refer to the same concept (see Fig. 3a). Specifically, we assess the correctness of $\widehat{y}$ using all four considered proprietary MLLMs and count it as correct if at least two models agree with this class (see details in Appendix E). In contrast to solving the BP task, which requires abstract reasoning abilities, the task of determining whether two answers refer to the same concept boils down to assessing the semantic similarity between two texts, and MLLMs are known to excel in such setup (Lu et al., 2023).
@@ 3.3 Binary Classification Formulations
To dive deeper into the AVR capabilities of the studied models, we cast the BP task into three binary classification settings, reducing the task’s difficulty.
Firstly, we provide the model with the image of the whole matrix ${\cal BP}^{X}$ along with a possible solution, and the model is prompted to generate a binary score assessing the correctness of the provided answer (Fig. 3b). Two settings are considered, in which the solution is formed by either the actual ground-truth answer (the expected answer is yes), or by an incorrect answer taken from a different matrix $\mathcal{BP}^{\cal X}$, where ${\cal X}=(X+20)\bmod 100$ (the expected answer is no). We manually reviewed the resulting pairings to confirm that the selected negative concept was distinct from the original concept. We refer to these two classification setups as Ground-Truth and Incorrect Label, resp.
@@
Secondly, we follow a setting exemplified in the Bongard-LOGO dataset (Nie et al., 2020) in which two test images have to be classified to different sides of the problem (Fig. 3c). To this end, we take two test images corresponding to the respective sides of the matrix, randomly shuffle the images, and request the model to determine the side to which each image belongs. In synthetic BPs we create the test set by removing the 6th image from each side of the matrix, while in BPs from Bongard HOI, Bongard-OpenWorld and Bongard-RWR we use the additional test images. We refer to this setup as Images to Sides.
Prompts for all three binary classification formulations are listed in Appendix I.3.
4 Bongard-RWR